# A Theoretical Perspective on Why Stochastic Population Update Needs an Archive in Evolutionary Multi-objective Optimization

### v2

@@ 4 Running Time Analysis of NSGA-II
We first introduce NGSA-II with SPU, and re-prove an upper bound on its expected running time for solving OneJumpZeroJump, which is much tighter than the previously known bound Bian et al. (2024b). Then, we prove that using an archive is helpful in this case, reducing the running time.
@@ 4.1 NSGA-II with SPU
Next, we introduce the stochastic population update (SPU) method for NSGA-II. During population updates, SPU randomly selects a fixed proportion $p_{s}$ of the current population and the offspring solutions to directly survive into the next generation and the removed part is selected from the rest solutions. This implies that each solution, including the worst solution in the population, has at least a probability $p_{s}$ of surviving to the next generation. Similar to the definition in Bian et al. (2024b), SPU of NSGA-II as presented in Algorithm 6 is used to replace the original Population Update procedure in line 14 of Algorithm 4. Note that $p_{s}$ is set to $1/4$ in Bian et al. (2024b), while we consider a general $p_{s}$ here.
@@
The expected running time of NSGA-II with SPU for solving the OneJumpZeroJump problem with $k>8e^{2}$ has been proven to be $O(\mu k(n/2)^{k})$ Bian et al. (2024b), which is better than that, i.e., $O(\mu n^{k})$, of the original NSGA-II Doerr and Qu (2023a). Here, we re-prove a tighter upper bound on the expected running time of NSGA-II with SPU for solving OneJumpZeroJump, as shown in Theorem 4. It is also more general, as it considers a survival probability $p_{s}\in[1/2\mu,1/2-o(1/\mu))$, rather than the specific probabilities $p_{s}=1/4$ analyzed in Bian et al. (2024b).
@@
Bian et al. Bian et al. (2024b) analyzed NSGA-II solving OneJumpZeroJump with $k>8e^{2}$, $p_{s}=1/4$, and $p_{c}=0$, deriving an expected running time of $O(\mu\sqrt{k}(n/2)^{k})$. In contrast, Theorem 4 gives a tighter bound $O(\mu n^{k}\cdot(e\ln(4e)/k)^{k-1})$, with an improvement ratio of $\Theta(\sqrt{k}(k/(2e\ln(4e)))^{k-1})$, which can be exponential when $k$ is large, e.g., $k=n/8$. Our analysis leads to improvement because, in the second phase of finding the extreme Pareto optimal solution $1^{n}$, we consider an optimal number $M+1$ of jumps across the gap of the dominated solution set (i.e., $\{\bm{x}\mid|\bm{x}|_{1}\in[n-k+1..n-1]\}$), rather than a continuous sequence of $k$ jumps from $|\bm{x}|_{1}=n-k$ to $n$ in Bian et al. (2024b).

### v3

@@ 4 Running Time Analysis of NSGA-II
We first introduce NGSA-II with SPU, and re-prove an upper bound on its expected running time for solving OneJumpZeroJump, which is much tighter than the previously known bound Bian et al. (2025). Then, we prove that using an archive is helpful in this case, reducing the running time.
@@ 4.1 NSGA-II with SPU
Next, we introduce the stochastic population update (SPU) method for NSGA-II. During population updates, SPU randomly selects a fixed proportion $p_{s}$ of the current population and the offspring solutions to directly survive into the next generation and the removed part is selected from the rest solutions. This implies that each solution, including the worst solution in the population, has at least a probability $p_{s}$ of surviving to the next generation. Similar to the definition in Bian et al. (2025), SPU of NSGA-II as presented in Algorithm 6 is used to replace the original Population Update procedure in line 14 of Algorithm 4. Note that $p_{s}$ is set to $1/4$ in Bian et al. (2025), while we consider a general $p_{s}$ here.
@@
The expected running time of NSGA-II with SPU for solving the OneJumpZeroJump problem with $k>8e^{2}$ has been proven to be $O(\mu k(n/2)^{k})$ Bian et al. (2025), which is better than that, i.e., $O(\mu n^{k})$, of the original NSGA-II Doerr and Qu (2023a). Here, we re-prove a tighter upper bound on the expected running time of NSGA-II with SPU for solving OneJumpZeroJump, as shown in Theorem 4. It is also more general, as it considers a survival probability $p_{s}\in[1/2\mu,1/2-o(1/\mu))$, rather than the specific probabilities $p_{s}=1/4$ analyzed in Bian et al. (2025).
@@
Bian et al. Bian et al. (2025) analyzed NSGA-II solving OneJumpZeroJump with $k>8e^{2}$, $p_{s}=1/4$, and $p_{c}=0$, deriving an expected running time of $O(\mu\sqrt{k}(n/2)^{k})$. In contrast, Theorem 4 gives a tighter bound $O(\mu n^{k}\cdot(e\ln(4e)/k)^{k-1})$, with an improvement ratio of $\Theta(\sqrt{k}(k/(2e\ln(4e)))^{k-1})$, which can be exponential when $k$ is large, e.g., $k=n/8$. Our analysis leads to improvement because, in the second phase of finding the extreme Pareto optimal solution $1^{n}$, we consider an optimal number $M+1$ of jumps across the gap of the dominated solution set (i.e., $\{\bm{x}\mid|\bm{x}|_{1}\in[n-k+1..n-1]\}$), rather than a continuous sequence of $k$ jumps from $|\bm{x}|_{1}=n-k$ to $n$ in Bian et al. (2025).